A father and his son are 50 years old and 20 years old respectively. Find the ratio of the father's age to his son' sage, when
a. The son was 10 years old
B. The father was 35 years old
C. The son will be 40 years old
D. The father will be 75 years old

Answer by lwsshak3(11628)

A father and his son are 50 years old and 20 years old respectively. Find the ratio of the father's age to his son' sage, when
a. The son was 10 years old
ratio:(50-10)/(20-10)=40/10=4
B. The father was 35 years old
ratio:(50-15)/(20-15)=35/5=7
C. The son will be 40 years old
ratio:(50+20)/(20+20)=70/40=7/4
D. The father will be 75 years old
ratio:(50+25)/(20+25)=75/45=5/3